# Maximal regularity for evolution equations with critical singular perturbations

### v1

@@ 3.3. Main result
By the maximal $L^{p}(0,T,w_{\kappa};X_{0})$-regularity and maximal $L^{r}(0,T,w_{\nu};X_{\gamma})$-regularity from Assumption 3.7, and by Lemma 2.2, we have a unique strong solution to (3.27). The claimed estimates follow from (2.6). ∎
@@
Let Assumptions 3.7 and 3.9 hold and let AA satisfy (3.18). Then for any $\tau\in(0,T]$, $u_{0}\in(X_{0},X_{1})_{1-\frac{1+\kappa}{p},p}$, $f\in{L_{+}}(0,\tau)$ and $v\in L^{p}(0,\tau,w_{\kappa};X_{1})$, there exists a unique strong solution $u\in\mathrm{MR}_{+}(0,\tau)$ to
@@
on $0,\tau]$. Furthermore, the following estimates hold:
$\|u\|_{\mathrm{MR}_{+}(0,\tau)}\leq$ $(M_{T}(1\vee C_{A})+1)\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}}+M_{T}\big(b_{\scriptscriptstyle[0,\tau]}\|v\|_{L^{p}(0,\tau,w_{\kappa};X_{1})}+\|f\|_{{L_{+}}(0,\tau)}\big),$
$\|u\|_{L^{p}(0,\tau,w_{\kappa};X_{1})}\leq$ $(M_{T}(1\vee C_{A})+1)\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}}$
(3.30) $\qquad+M_{T}K\big(b_{\scriptscriptstyle[0,\tau]}\|v\|_{L^{p}(0,\tau,w_{\kappa};X_{1})}+\|f\|_{{L_{+}}(0,\tau)}\big),$
where $M_{T}$ is defined by (3.16), $C_{A}$ is the constant from (3.18) and $K$ is a constant independent of $\tau,T,u_{0},f$, and $v$.
@@
Let $\tau\in(0,T]$. First, consider the problem
(3.31) $\begin{cases}\tilde{u}^{\prime}+A\tilde{u}=0,\\ \tilde{u}(0)=u_{0}.\end{cases}$
Observe that $L^{p}(0,T,w_{\kappa};X_{1})\cap W^{1,p}(0,T,w_{\kappa};X_{0})\hookrightarrow L^{p}(0,T,w_{\kappa};X_{1})$ non-expansively, thus (3.18) implies that (2.5) holds with the same constant $C_{A}$. Combined with the maximal regularity in Assumption 3.7(2) and Lemmas 2.2 and 2.3, it follows that there exists a unique strong solution $\tilde{u}\in L^{p}(0,\tau,w_{\kappa};X_{1})\cap W^{1,p}(0,\tau,w_{\kappa};X_{0})$ to (3.31) and
(3.32) $\|\tilde{u}\|_{L^{p}(0,\tau,w_{\kappa};X_{1})\cap W^{1,p}(0,\tau,w_{\kappa};X_{0})}\leq(M_{T}(1\vee C_{A})+1)\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}}.$
Next, let $f=f_{0}+\sum_{i=1}^{k}f_{i}$ with $f_{0}\in L^{p}(0,\tau,w_{\kappa};X_{0})$ and $f_{i}\in L^{r_{i}}(0,\tau,w_{\nu_{i}};X_{\gamma_{i}})$. Put $g_{0}=f_{0}$ and $g_{i}=f_{i}-B_{i}v$ for $1\leq i\leq k$. For $i=0$, set $(r_{0},\nu_{0},\gamma_{0})=(p,\kappa,0)$. By Lemma 3.16, we have $g_{i}\in L^{r_{i}}(0,\tau,w_{\nu_{i}};X_{\gamma_{i}})$ for $1\leq i\leq k$. By the maximal regularity in Assumption 3.7(3), for $0\leq i\leq k$, there exists a unique strong solution $u^{i}\in L^{r_{i}}(0,\tau,w_{\nu_{i}};X_{1+\gamma_{i}})\cap W^{1,r_{i}}(0,\tau,w_{\nu_{i}};X_{\gamma_{i}})$ to
@@
It follows that $u\coloneqq\tilde{u}+\sum_{i=0}^{k}u^{i}$ is a strong solution to (3.29).
@@
It remains to prove the energy estimate. Let $f\in{L_{+}}(0,\tau)$ and let $f=f_{0}+\sum_{i=1}^{k}f_{i}$ be an arbitrary composition for $f$ with $f_{i}\in L^{r_{i}}(0,\tau,w_{\nu_{i}};X_{\gamma_{i}})$. We can define $\tilde{u}$ and $u^{i}$ as above (now with potentially different $f_{i}$) and obtain that $\tilde{u}+\sum_{i=0}^{k}u^{i}$ is a strong solution to (3.29) on $0,\tau]$ that belongs to $\mathrm{MR}_{+}(0,\tau)$, thus uniqueness gives $u=\tilde{u}+\sum_{i=0}^{k}u^{i}$. Moreover, by (3.32) and Lemma 2.2 applied to $u^{i}$, we have for all $\tau\in(0,T]$:
@@
(3.34) $\begin{split}&\leq(M_{T}(1\vee C_{A})+1)\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}}+\textstyle{\sum_{i=0}^{k}}M_{T}\|g_{i}\|_{L^{r_{i}}(0,\tau,w_{\nu_{i}};X_{\gamma_{i}})}\\ &\leq(M_{T}(1\vee C_{A})+1)\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}}+M_{T}b_{\scriptscriptstyle[0,\tau]}\|v\|_{L^{p}(0,\tau,w_{\kappa};X_{1})}\\ &\qquad\qquad+M_{T}\|f_{0}\|_{L^{p}(0,\tau,w_{\kappa};X_{0})}+M_{T}\textstyle{\sum_{i=1}^{k}}\|f_{i}\|_{L^{r_{i}}(0,\tau,w_{\nu_{i}};X_{\gamma_{i}})},\end{split}$
where we applied (3.32) for the initial value part and combined Lemma 3.15 with Lemma 3.16 for the $g_{i}$-terms. Taking the infimum over all decompositions $f_{0}+\sum_{i=1}^{k}f_{i}$ of $f$, we obtain the $\mathrm{MR}_{+}$-estimate.
Finally, we prove the $L^{p}$-estimate. For $i=1,\ldots,k$, Corollary 3.6 and $u^{i}(0)=0$ imply that $\|u^{i}\|_{L^{p}(0,\tau,w_{\kappa};X_{1})}\leq K\|u^{i}\|_{\mathrm{MR}_{+}(0,\tau)}$ for a $\tau$-independent constant $K$. Furthermore, (3.32) gives $\|\tilde{u}\|_{L^{p}(0,\tau,w_{\kappa};X_{1})}\leq(M_{T}(1\vee C_{A})+1)\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}}$. Combining this with (3.34) yields (3.30). ∎
@@
where we let $C_{0}\coloneqq\max\{(M_{T}(1\vee C_{A})+1),M_{T}K\}$, with $M_{T}$, $C_{A}$ and $K$ the constants from Proposition 3.17, and we use that $\tau\in(0,T]$. Note that for (3.35), the constant $M_{T}K$ suffices, but we have defined $C_{0}$ as a maximum for later use.
@@
Combined with (3.30) for $v=0$, we obtain
@@
Moreover, we have $f{{\bf 1}}_{(\tau_{n},\tau_{n+1})}-B\tilde{w}{{\bf 1}}_{(\tau_{n},\tau_{n+1})}\in{L_{+}}(0,T)$ by Lemma 3.16. Therefore, for any $v\in L^{p}(0,\tau_{n+1},w_{\kappa};X_{1})$, Proposition 3.17 yields a unique strong solution $w\eqqcolon\Phi_{n}(v)$ to (3.40) with $w\in\mathrm{MR}_{+}(0,\tau_{n+1})\subset L^{p}(0,\tau_{n+1},w_{\kappa};X_{1})$. Moreover, by virtue of (3.30) and (3.38), the analog of (3.36) holds:
@@
The claimed estimate (3.19) follows by combining (3.45) from Step 3 with an application of Proposition 3.17 with $u=v$. The estimate (3.20) then follows from (3.19) and Corollary 3.6. ∎
@@
$\mathrm{MR}_{\rm tr}(0,T)\coloneqq L^{p}(0,T,w_{\kappa};X_{1})\cap C([0,T];X_{1-\frac{1+\kappa}{p},p})\cap C_{\nicefrac{{\kappa}}{{p}}}([0,T];X_{1-\frac{1}{p},p}),$
@@
3.5. Further embeddings of the maximal regularity space
We recall that in (3.20) we already stated continuity properties of solutions, which were derived from Corollary 3.6. In this subsection, we explore further regularity properties, in particular, fractional smoothness in time.
We now establish useful embeddings for the space $\mathrm{MR}_{+}(0,T)$ (see (3.6)). Note that all embeddings in this subsection rely on an extension operator. Consequently, the operator norm of each embedding depends on $T$ and tends to infinity as $T\downarrow 0$, unless the domain is restricted to functions in $\mathrm{MR}_{+}(0,T)$ vanishing at $t=0$.

### v2

@@ 3.3. Main result
(i): By the maximal $L^{p}(0,T,w_{\kappa};X_{0})$-regularity from Assumption 3.7(2) and by Lemma 2.2, we have a unique strong solution in $L^{p}(0,\tau,w_{\kappa};X_{1})\cap W^{1,p}(0,\tau,w_{\kappa};X_{0})$ to (3.25). The claimed estimate follows from (2.6).
(ii): By the maximal $L^{r}(0,T,w_{\nu};X_{\gamma})$-regularity in Assumption 3.7(3), there exists a unique strong solution $v\in L^{r}(0,\tau,w_{\nu};X_{1+\gamma})\cap W^{1,r}(0,\tau,w_{\nu};X_{\gamma})$ to $v^{\prime}+A_{r,\nu,\gamma}v=f$, $v(0)=0$. Moreover, defining $F(0,\tau)$ as in the statement, using a bounded extension operator $\mathrm{E}\colon F(0,\tau)\to F(0,T)$ (see [18, Prop. L.4.5]) and using (3.16), we have
$\|(A_{r,\nu,\gamma}-A)v\|_{L^{p}(0,\tau,w_{\kappa};X_{0})}$ $\leq\|(A_{r,\nu,\gamma}-A)\mathrm{E}v\|_{L^{p}(0,T,w_{\kappa};X_{0})}$
(3.26) $\leq(C_{A_{r,\nu,\gamma}}+C_{A})K\|\mathrm{E}\|\|v\|_{L^{r}(0,\tau,w_{\nu};X_{\gamma+1})\cap W^{1,r}(0,\tau,w_{\nu};X_{\gamma})},$
where $K$ is the embedding constant of Corollary 3.6, which is $\tau$- and $T$-independent since $v(0)=0$, and $\|\mathrm{E}\|$ depends only on $p,\kappa,r,$ and $\nu$ (see [18, Prop. L.4.5(4)]). Therefore, by Assumption 3.7(2), we have a unique strong solution $w\in L^{p}(0,\tau,w_{\kappa};X_{1})\cap W^{1,p}(0,\tau,w_{\kappa};X_{0})$ to $w^{\prime}+Aw=(A_{r,\nu,\gamma}-A)v$, $w(0)=0$.
Now put $u\coloneqq v+w$. Then $u\in F(0,\tau)$ and $u$ is a strong solution to (3.25). Moreover, by maximal regularity, Lemma 2.2, and (3.4),
$\|u\|_{F(0,\tau)}$ $\leq\|v\|_{L^{r}(0,\tau,w_{\nu};X_{\gamma+1})\cap W^{1,r}(0,\tau,w_{\nu};X_{\gamma})}+\|w\|_{L^{p}(0,\tau,w_{\kappa};X_{1})\cap W^{1,p}(0,\tau,w_{\kappa};X_{0})}$
$\leq\big(1+M_{p,\kappa,A}(0,T)(C_{A_{r,\nu,\gamma}}+C_{A})K\|\mathrm{E}\|\big)\|v\|_{L^{r}(0,\tau,w_{\nu};X_{\gamma+1})\cap W^{1,r}(0,\tau,w_{\nu};X_{\gamma})}$
$\leq M_{r,\nu,A_{r,\nu,\gamma}}(0,T)\big(1+M_{p,\kappa,A}(0,T)(C_{A_{r,\nu,\gamma}}+C_{A})K\|\mathrm{E}\|\big)\|f\|_{L^{r}(0,\tau,w_{\nu};X_{\gamma})}.$
The following lemma is an extension of Lemma 2.3, in which we assume that $A$ has the mapping property (3.16) instead of (2.5). We consider the homogeneous problem
(3.27) $\begin{cases}{u}^{\prime}+A{u}=0,\\ {u}(0)=u_{0}.\end{cases}$
Lemma 3.15.
Let Assumption 1.1 be satisfied. Let $p\in(1,\infty)$ and $\kappa\in[0,p-1)$. Suppose that the condition for $A$ in (3.16) holds and $A$ has maximal $L^{p}(0,T,w_{\kappa};X_{0})$-regularity.
For every $\tau\in(0,T]$ and $u_{0}\in X_{1-\frac{1+\kappa}{p},p}$, there exists a unique strong solution $u\in L^{p}(0,\tau,w_{\kappa};X_{1})\cap W^{1,p}(0,\tau,w_{\kappa};X_{0})$ to (3.27), and
$\|u\|_{L^{p}(0,\tau,w_{\kappa};X_{1})\cap W^{1,p}(0,\tau,w_{\kappa};X_{0})}+\|u\|_{C([0,\tau];X_{1-\frac{1+\kappa}{p},p})}+\|u\|_{C_{\nicefrac{{\kappa}}{{p}}}((0,\tau];X_{1-\frac{1}{p},p})}\leq R_{T}\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}},$
where $R_{T}$ is a constant that depends only on $p,\kappa$, and nondecreasingly on $M_{p,\kappa,A}(0,T)$ and $C_{A}$. In particular, $R_{T}$ is independent of $\tau$ and $u_{0}$.
Proof.
Uniqueness follows from (2.6). For the existence and estimate, we exploit the trace method for real interpolation (see [18, §L.2] and [29, §1.8]). Using the definition of the $\mathrm{MR}^{p}$-space in (2.2), let $z\in\mathrm{MR}^{p}(0,\infty,w_{\kappa})$ be such that $z(0)=u_{0}$. Note that $z^{\prime},Az\in L^{p}(0,\tau,w_{\kappa};X_{0})$ thanks to (3.16), (2.3) and (2.4). Let $v\in\mathrm{MR}^{p}(0,\tau,w_{\kappa})$ be the unique solution to the problem $v^{\prime}+Av=-z^{\prime}-Az$ with $v(0)=0$. Then, $u\coloneqq v+z$ satisfies (3.27). It remains to prove the estimates for $u$.
Write $M\coloneqq M_{p,\kappa,A}(0,T)$. By maximal regularity and by (3.16), (2.3) and (2.4), we have
$\|v\|_{\mathrm{MR}^{p}(0,\tau,w_{\kappa})}$ $\leq M\big(\|z^{\prime}\|_{L^{p}(0,T,w_{\kappa};X_{0})}+\|Az\|_{L^{p}(0,T,w_{\kappa};X_{0})}\big)$
$\leq M\|z^{\prime}\|_{L^{p}(0,T,w_{\kappa};X_{0})}$
$\quad+MC_{A}\big(\|z\|_{L^{p}(0,T,w_{\kappa};X_{1})}+\|z\|_{C([0,T];X_{1-\frac{1+\kappa}{p},p})}+\|z\|_{C_{\nicefrac{{\kappa}}{{p}}}((0,T];X_{1-\frac{1}{p},p})}\big)$
$\leq M(1\vee C_{A})\|z\|_{\mathrm{MR}^{p}(0,\infty,w_{\kappa})}+MC_{A}\big(\|z\|_{C([0,\infty);X_{1-\frac{1+\kappa}{p},p})}+\|z\|_{C_{\nicefrac{{\kappa}}{{p}}}((0,\infty);X_{1-\frac{1}{p},p})}\big)$
(3.28) $\leq\big(M(1\vee C_{A})+MC_{A}(K_{\kappa,p}+K_{0,p})\big)\|z\|_{\mathrm{MR}^{p}(0,\infty,w_{\kappa})},$
where $K_{\kappa,p}\coloneqq 64p(\frac{1}{p-1+\kappa}\vee\frac{1}{1+\kappa})$ and $K_{0,p}\coloneqq 64p(\frac{1}{p-1}\vee 1)$ are constants independent of $\tau$ and $T$, since (2.3) and (2.4) are applied on $0,\infty)$ in the last line (see [18, Th. L.4.1] for the constants).
We conclude that
$\|u\|_{\mathrm{MR}^{p}(0,\tau,w_{\kappa})}$ $\leq\|v\|_{\mathrm{MR}^{p}(0,\tau,w_{\kappa})}+\|z\|_{\mathrm{MR}^{p}(0,\infty,w_{\kappa})}$
$\leq\big(M(1\vee C_{A})+1+MC_{A}(K_{\kappa,p}+K_{0,p})\big)\|z\|_{\mathrm{MR}^{p}(0,\infty,w_{\kappa})}.$
Taking the infimum over all $z\in\mathrm{MR}^{p}(0,\infty,w_{\kappa})$, the first stated estimate follows from the trace method for real interpolation (with an additional constant $\frac{12p}{1+\kappa}$, see [18, Th. L.2.3]).
Moreover, by (2.3),
$\|u\|_{C([0,\tau];X_{1-\frac{1+\kappa}{p},p})}$ $\leq\|v\|_{C([0,\tau];X_{1-\frac{1+\kappa}{p},p})}+\|z\|_{C([0,\infty);X_{1-\frac{1+\kappa}{p},p})}$
$\leq K_{\kappa,p}C_{\kappa,p}\|v\|_{\mathrm{MR}^{p}(0,\tau,w_{\kappa})}+K_{\kappa,p}\|z\|_{\mathrm{MR}^{p}(0,\infty,w_{\kappa})},$
where $C_{\kappa,p}\coloneqq 2+3(1-\frac{\kappa}{p-1})^{\frac{p-1}{p}}$ is a constant that is also independent of $\tau$ and $T$ since $v(0)=0$, see [18, Th. L.4.1, Prop. L.4.5].
Similarly, by (2.4),
$\|u\|_{C_{\nicefrac{{\kappa}}{{p}}}((0,\tau];X_{1-\frac{1}{p},p})}$ $\leq\|v\|_{C_{\nicefrac{{\kappa}}{{p}}}((0,\tau];X_{1-\frac{1}{p},p})}+\|z\|_{C_{\nicefrac{{\kappa}}{{p}}}((0,\infty);X_{1-\frac{1}{p},p})}$
$\leq K_{0,p}C_{\kappa,p}\|v\|_{\mathrm{MR}^{p}(0,\tau,w_{\kappa})}+K_{0,p}\|z\|_{\mathrm{MR}^{p}(0,\infty,w_{\kappa})},$
see again [18, Th. L.4.1, Prop. L.4.5]) and recall that $v(0)=0$. Combining these estimates with (3.28), and taking the infimum over all $z\in\mathrm{MR}^{p}(0,\infty,w_{\kappa})$, the trace method for real interpolation completes the proof of the last two estimates (again with extra constant $\frac{12p}{1+\kappa}$, see [18, Th. L.2.3]). ∎
@@
Let Assumptions 3.7 and 3.9 hold. Then for any $\tau\in(0,T]$, $u_{0}\in(X_{0},X_{1})_{1-\frac{1+\kappa}{p},p}$, $f\in{L_{+}}(0,\tau)$ and $v\in L^{p}(0,\tau,w_{\kappa};X_{1})$, there exists a unique strong solution $u\in\mathrm{MR}_{+}(0,\tau)$ to
@@
on $0,\tau]$. Furthermore, the following estimate holds:
$\|u\|_{\mathrm{MR}_{+}(0,\tau)}+\|u\|_{L^{p}(0,\tau,w_{\kappa};X_{1})}+\|u\|_{C([0,\tau];X_{1-\frac{1+\kappa}{p},p})}+\|u\|_{C_{\nicefrac{{\kappa}}{{p}}}((0,\tau];X_{1-\frac{1}{p},p})}$
$\qquad\leq R_{T}\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}}+L_{T}\big(b_{\scriptscriptstyle[0,\tau]}\|v\|_{L^{p}(0,\tau,w_{\kappa};X_{1})}+\|f\|_{{L_{+}}(0,\tau)}\big),$
where $R_{T}$ is the constant from Lemma 3.15, and $L_{T}$ is a constant depending only on $p,\kappa$, and $r_{i},\nu_{i},\gamma_{i}$ for $1\leq i\leq k$, and nondecreasingly on $M_{T}$ defined by (3.18), and $C_{A}$ and $C_{A_{r_{i},\nu_{i},\gamma_{i}}}$ from (3.16). In particular, $R_{T}$ and $L_{T}$ are independent of $\tau,u_{0},f$, and $v$.
@@
Let $\tau\in(0,T]$. First, by Lemma 3.15, the problem $\tilde{u}^{\prime}+A\tilde{u}=0,\tilde{u}(0)=u_{0}$ has a unique strong solution $\tilde{u}\in L^{p}(0,\tau,w_{\kappa};X_{1})\cap W^{1,p}(0,\tau,w_{\kappa};X_{0})$, and
(3.31) $\|\tilde{u}\|_{L^{p}(0,\tau,w_{\kappa};X_{1})\cap W^{1,p}(0,\tau,w_{\kappa};X_{0})}+\|u\|_{C([0,\tau];X_{1-\frac{1+\kappa}{p},p})}+\|u\|_{C_{\nicefrac{{\kappa}}{{p}}}((0,\tau];X_{1-\frac{1}{p},p})}\leq R_{T}\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}},$
where $R_{T}$ is a constant that depends only on $p,\kappa$, and nondecreasingly on $M_{p,\kappa,A}(0,T)$ and $C_{A}$.
Next, let $f=f_{0}+\sum_{i=1}^{k}f_{i}$ with $f_{0}\in L^{p}(0,\tau,w_{\kappa};X_{0})$ and $f_{i}\in L^{r_{i}}(0,\tau,w_{\nu_{i}};X_{\gamma_{i}})$. Put $g_{0}=f_{0}$ and $g_{i}=f_{i}-B_{i}v$ for $1\leq i\leq k$. For $i=0$, set $(r_{0},\nu_{0},\gamma_{0})=(p,\kappa,0)$. By Lemma 3.16, we have $g_{i}\in L^{r_{i}}(0,\tau,w_{\nu_{i}};X_{\gamma_{i}})$ for $1\leq i\leq k$. By Lemma 3.14, there exists a strong solution $u^{i}\in\mathrm{MR}_{+}(0,\tau)$ to
@@
It follows that $u\coloneqq\tilde{u}+\sum_{i=0}^{k}u^{i}\in\mathrm{MR}_{+}(0,\tau)$ is a strong solution to (3.30).
@@
It remains to prove the energy estimate. Let $f\in{L_{+}}(0,\tau)$ and let $f=f_{0}+\sum_{i=1}^{k}f_{i}$ be an arbitrary composition for $f$ with $f_{i}\in L^{r_{i}}(0,\tau,w_{\nu_{i}};X_{\gamma_{i}})$. We can define $\tilde{u}$ and $u^{i}$ from Lemma 3.14 as above (now with potentially different $f_{i}$) and obtain that $\tilde{u}+\sum_{i=0}^{k}u^{i}$ is a strong solution to (3.30) on $0,\tau]$ that belongs to $\mathrm{MR}_{+}(0,\tau)$, thus uniqueness gives $u=\tilde{u}+\sum_{i=0}^{k}u^{i}$. Moreover, by (3.31) and by Lemmas 3.14 and 3.16, putting $C_{*}\coloneqq C_{A}+\max_{1\leq i\leq k}C_{A_{\gamma_{i},\nu_{i},r_{i}}}$, we have
@@
$\leq R_{T}\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}}+M_{T}\|f_{0}\|_{L^{p}(0,\tau,w_{\kappa};X_{0})}+\textstyle{\sum_{i=1}^{k}}M_{T}(1+M_{T}LC_{*})\|g_{i}\|_{L^{r_{i}}(0,\tau,w_{\nu_{i}};X_{\gamma_{i}})}$
$\leq R_{T}\|u_{0}\|_{X_{1-\frac{1+\kappa}{p},p}}+M_{T}\|f_{0}\|_{L^{p}(0,\tau,w_{\kappa};X_{0})}$
(3.33) $\qquad\qquad+M_{T}(1+M_{T}LC_{*})\Big(b_{\scriptscriptstyle[0,\tau]}\|v\|_{L^{p}(0,\tau,w_{\kappa};X_{1})}+\textstyle{\sum_{i=1}^{k}}\|f_{i}\|_{L^{r_{i}}(0,\tau,w_{\nu_{i}};X_{\gamma_{i}})}\Big),$
where we recall (3.17) for the last line. Taking the infimum over all decompositions $f_{0}+\sum_{i=1}^{k}f_{i}$ of $f$, we obtain the $\mathrm{MR}_{+}$-estimate.
Finally, we prove the remaining estimates. For $i=1,\ldots,k$, Corollary 3.6 and $u^{i}(0)=0$ imply
$\|u^{i}\|_{L^{p}(0,\tau,w_{\kappa};X_{1})}+\|u^{i}\|_{C([0,\tau];X_{1-\frac{1+\kappa}{p},p})}+\|u^{i}\|_{C_{\nicefrac{{\kappa}}{{p}}}((0,\tau];X_{1-\frac{1}{p},p})}\leq K\|u^{i}\|_{\mathrm{MR}_{+}(0,\tau)}$
for a constant $K$ depending only on $p,\kappa,$ and $r_{i},\nu_{i},\gamma_{i}$ for $1\leq i\leq k$. Moreover, (3.31) already provided the necessary estimates for $\tilde{u}$. Combining these estimates with those in (3.33) completes the proof. ∎
@@
where we let $C_{0}\coloneqq R_{T}\vee L_{T}$, with $R_{T}$ and $L_{T}$ the constants from Proposition 3.17, and we use that $\tau\in(0,T]$. Note that for (3.34), the constant $L_{T}$ suffices, but we define $C_{0}$ as the maximum for later use.
@@
Combined with Proposition 3.17 applied to $v=0$, we obtain
@@
Moreover, we have $f{{\bf 1}}_{(\tau_{n},\tau_{n+1})}-B\tilde{w}{{\bf 1}}_{(\tau_{n},\tau_{n+1})}\in{L_{+}}(0,T)$ by Lemma 3.16. Therefore, for any $v\in L^{p}(0,\tau_{n+1},w_{\kappa};X_{1})$, Proposition 3.17 yields a unique strong solution $w\eqqcolon\Phi_{n}(v)$ to (3.39) with $w\in\mathrm{MR}_{+}(0,\tau_{n+1})\subset L^{p}(0,\tau_{n+1},w_{\kappa};X_{1})$. Moreover, by virtue of Proposition 3.17 and (3.37), the analog of (3.35) holds:
@@
The estimate in the statement of Theorem 3.11 follows by combining (3.44) from Step 3 with an application of Proposition 3.17 with $u=v$. ∎
Remark 3.19.
Regarding the mapping condition (3.16) in Assumption 3.7, there is some flexibility in choosing the function space for u without invalidating the results of Theorem 3.11. Indeed, in (3.16), it suffices to estimate by $C_{A}\|u\|_{E(0,T)}$ for any normed space $E(0,T)$ such that $\mathrm{MR}_{+}(0,T)\hookrightarrow E(0,T)\hookrightarrow L^{p}(0,T,w_{\kappa};X_{1})$, as seen from the proof of Proposition 3.17. However, choosing the smallest space $E(0,T)=\mathrm{MR}_{+}(0,T)$ can lead to a constant $C_{A}$ that diverges as $T\downarrow 0$, which may be undesirable. Our current choice prevents this behavior and covers virtually all cases of interest.
@@
$\mathrm{MR}_{\rm tr}(0,T)\coloneqq L^{p}(0,T,w_{\kappa};X_{1})\cap C([0,T];X_{1-\frac{1+\kappa}{p},p})\cap C_{\nicefrac{{\kappa}}{{p}}}((0,T];X_{1-\frac{1}{p},p}),$
@@
Remark 3.21 (Generalization to intermediate spaces).
Theorem 3.11 can be extended to perturbations $B_{i}\colon[0,T]\to\mathcal{L}(Y_{\beta_{i}},X_{\gamma_{i}})$ acting on intermediate spaces. For $i=1,\dots,k$, assume that $\beta_{i}\in[1-\frac{1+\kappa}{p},1]$ and $Y_{\beta_{i}}$ is a Banach space satisfying $X_{\beta_{i},1}\hookrightarrow Y_{\beta_{i}}$. We replace the integrability condition in Assumption 3.9 by $\|B_{i}(\cdot)\|_{\mathcal{L}(Y_{\beta_{i}},X_{\gamma_{i}})}\leq b_{i}(\cdot)$, where $b_{i}\in L^{q_{i}}(0,T,w_{\mu_{i}})$ with parameters
$q_{i}\coloneqq\frac{pr_{i}}{p-r_{i}\theta_{i}},\quad\mu_{i}\coloneqq\frac{\nu_{i}p-\kappa r_{i}\theta_{i}}{p-r_{i}\theta_{i}},\qquad\text{where }\,\theta_{i}\coloneqq 1-\frac{(1-\beta_{i})p}{1+\kappa}.$
The well-posedness in $\mathrm{MR}_{+}(0,T)$ and estimates of Theorem 3.11 remain valid in this setting. The proof follows the same fixed-point strategy but relies on the auxiliary space $E_{p,\kappa}(0,\tau)\coloneqq L^{p}(0,\tau,w_{\kappa};X_{1})\cap L^{\infty}(0,\tau;X_{1-\frac{1+\kappa}{p},p})$. Crucially, using the reiteration theorem and Assumption 3.7(1), one has $X_{\beta_{i},1}=(X_{1-\frac{1+\kappa}{p},p},X_{1})_{\theta_{i},1}$, which implies the following interpolation estimate:
$\|B_{i}v\|_{L^{r_{i}}(0,\tau,w_{\nu_{i}};X_{\gamma_{i}})}\leq C\|b_{i}\|_{L^{q_{i}}(0,\tau,w_{\mu_{i}})}\|v\|_{L^{\infty}(0,\tau;X_{1-\frac{1+\kappa}{p},p})}^{1-\theta_{i}}\|v\|_{L^{p}(0,\tau,w_{\kappa};X_{1})}^{\theta_{i}},\qquad v\in E_{p,\kappa}(0,\tau).$
This allows one to reprove Proposition 3.17 for $v\in E_{p,\kappa}(0,\tau)$ (with $\|v\|_{E_{p,\kappa}(0,\tau)}$ in the estimate) and close the contraction arguments in the proof of Theorem 3.11 in the spaces $E_{p,\kappa}(0,\tau)$ (Step 1) and $E_{p,\kappa}(0,\tau_{n+1})$ (Step 3).
3.5. Further embeddings of the mixed-scale maximal regularity space
We recall that Theorem 3.11 already provided continuity properties of solutions, which were derived from Corollary 3.6. In this subsection, we explore further regularity properties, in particular, fractional smoothness in time.
We establish useful embeddings for the space $\mathrm{MR}_{+}(0,T)$ (see (3.6)). Note that all embeddings in this subsection rely on an extension operator. Consequently, the operator norm of each embedding depends on $T$ and tends to infinity as $T\downarrow 0$, unless the domain is restricted to functions in $\mathrm{MR}_{+}(0,T)$ vanishing at $t=0$.